OpenStudy (anonymous):

Simplify 10 √150 -√20 +√45

OpenStudy (anonymous):

50 √6 -2√5 +3√5=50 √6 +√5

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

what can u simplify radical 150 to

OpenStudy (anonymous):

try to devise 150 by the biggest square 25 and 9 and 4

OpenStudy (anonymous):

okay I got that than you simplify radical 30 by radical 4 and radical 5 and u keep the minus sign

OpenStudy (anonymous):

right

OpenStudy (anonymous):

10 √150=10 √25*6 =5*10 √6

OpenStudy (anonymous):

Ya I got that

OpenStudy (anonymous):

when you break down radical 45 to radical 5 and radical 9